Jen wants to accumulate 400,000 at the end of 16 years. She deposits CX+10,000) into an investment account at the end of each of the first 8 years. She then deposits X into the same account at the end of each of the final 8 years. The interest rate on the account is a 6% nominal rate compounded annually. Find X 7.

8. Keri was offered a choice of two payment options to settle her claims in a car accident case. The first option would pay her a single lump payment of 1,500,000 immediately, which she would deposit into an account earning an effective annual interest rate of i. The second option would pay her 4 annual payments of 300,000 with the first payment being made immediately and the final one at time 3. Keri can invest the payments from the second option into an account crediting an effective annual interest rate of 10%. After 10 years, the account values of the two options are equal. Calculate i for the first option.

media%2Fd65%2Fd65aee42-4819-4a3c-ab04-e7

10. Alexa deposits $10,000 annually at the beginning of the year for 6 years into a special MSOE test account. The deposits earn 5% annually, and interest is paid out at the end of each year and reinvested at 400 into a second MSOE test account. Find the total value in the two test accounts at the end of 6 years.
Answer #1

Since you have posted multiple unrelated questions, i will answer the first one.

Question No. 7

Let's look at first investment made at the end of year 1. So this investment has N = 16 - 1 = 15 years to grow until maturity (i.e end of year 16). Interest rate, I = 6%.

So the FV factor associated with this investment = FV factor = (1 + I)N = (1 + 6%)15 = 2.3966

So, the future value of investment made at the end of year 1 = 2.3966 x amount invested = 2.3966 x (X + 10,000). We will now extend the same logic for investments made subsequently. We will thus create the following FV factor table:

End of year, P

Number of years until maturity

FV factor

P

N = 16 - P

FV = (1+6%)N

1

15

2.3966

2

14

2.2609

3

13

2.1329

4

12

2.0122

5

11

1.8983

6

10

1.7908

7

9

1.6895

8

8

1.5938

Subtotal for P = 1 to 8

15.7751

End of year, P

Number of years until maturity

FV factor

P

N = 16 - P

FV = (1+6%)N

9

7

1.5036

10

6

1.4185

11

5

1.3382

12

4

1.2625

13

3

1.1910

14

2

1.1236

15

1

1.0600

16

0

1.0000

Sub total for P = 9 to 16

9.8975

We have split the table in two parts because amount invested per year is different for the two blocks of time:

  • For P =1 to P =8; amount invested per annum = X + 10,000; so the future value of investments = Amount invested per annum x Sum of FV factors corresponding to the block of first 8 years = (X + 10,000) x 15.7751; Please see the table above for calculation of 15.7751
  • For P =9 to P =16; amount invested per annum = X; so the future value of investments = Amount invested per annum x Sum of FV factors corresponding to the block of next 8 years, starting from year 9 to year 16 = X x 9.8975; Please see the table above for calculation of 9.8975

So, the future value of these investments should add up to the target value of 400,000.

Hence, the equation in hand is:

(X + 10,000) x 15.7751 + X x 9.8975 = 400,000

Hence, 25.6725X + 157,751 = 400,000

Hence, X = 9,436.13